Find the heat in jouls produced in an electronic cord of resistance 30 ohms when a current of 5 ampears flow through it for 30 minutes
1
Expert's answer
2019-12-18T11:28:18-0500

"Q=I^2Rt"

"Q=(5)^2(30)(30)(60)=1350000\ J=1.35\ MJ"